a person is paid $153 for 9 baskets of dried flowers. To earn $450, how many baskets of dried flowers must the person produce?
Answer by farah(18)

step 1 - 153/9 = 17
step 2 - 450/17 = 26/8/17 or 26.47
answer : 26.47